# Detect, Investigate, Judge and Determine: A Knowledge-guided Framework for Few-shot Fake News Detection

### v3

@@ 5.2 Experimental Result
The main results, presented in Table 2, indicate that our proposed DKFND model surpasses all baselines across various metrics, encompassing traditional, LLM-based and hybrid methods. This underscores the effectiveness of our design and the advantages of enhancing the LLM through both inside and outside perspectives. Furthermore, several notable phenomena emerge from these results:
@@
Firstly, for most baselines and our DKFND model, the performance on the PolitiFact dataset exceeds that on the Gossipcop dataset, suggesting that Gossipcop presents greater difficulty. Specifically, PolitiFact consists of political news while Gossipcop pertains to the entertainment domain. This disparity is reasonable as political news typically exhibits more organized format and content, which facilitates the fake news detection process. Secondly, with the increase of training instances ($K$), most traditional fake news detection methods (e.g., ④ PSM) and hybrid methods ( 11 ARG) show improved performance. This is logical as more data enables better training of a supervised model, mitigating the lack of prior knowledge. However, an exception is observed in ① PROPANEWS, whose performance appears relatively unaffected by $K$. As introduced in Section 5.1.2, different from other methods, PROPANEWS designs a data constructing strategy to supplement original training set. This significantly offsets the impact of training data quantity. Moreover, for most LLM-based methods (⑦ $\sim$ 10), as outlined in Section 5.1.2, due to the limitation of maximum tokens, we all randomly select 6 samples as demonstrations. Hence, increasing the number of training instances does not substantially benefit the in-context learning of LLMs. Thirdly, although our DKFND model also faces the constraint of maximum token limitations, the $k$NN retrieval mechanism and inside judge designs enhance the utilization of increased training data, thereby achieving a certain degree of improvements with higher $K$ values. Fourthly, the hybrid method ( 11 ARG), benefiting from the joint modeling of traditional models and LLMs, obtains competitive performance. And compared to that, DKFND still maintains a significant advantage, particularly in scenarios with scarcer data. These observations further demonstrate the effectiveness of our designs from multiple perspectives.
5.3 Component Effectiveness Analysis
@@
In this subsection, we conduct ablation experiments to assess the effectiveness of different components within our model. Specifically, we first simplify the Integrated Determination part (Section 4.4.1), removing the detailed explanation in promot $P_{FD}$ (Eq.(10)). On the basis of this, we design ablated variants from inside and outside perspectives.
@@
  • DKFND-Inside: simplifying the design of Section 4.2.1 and 4.3.1 by randomly selecting samples from training set as demonstrations.
@@
  • DKFND-Outside: simplifying the design of Section 4.3.2 by randomly selecting retrieved outside information.
@@
Furthermore, considering that the backbone of DKFND is ⑦ Zephyr (Tunstall et al., 2023), we can jointly compare the results in Figure 3 and the “⑦ Zephyr” line in Table 2. This comparison further reveals that all three ablation variants outperform Zephyr, demonstrating that even in its simplified forms, leveraging both internal and external perspectives for fake news detection yields significant performance gains. This finding further validates the motivation behind our dual-perspective design.
5.4 Case Study
@@
More experimental analyses, such as Bad Case Analysis, can be found in Appendix AC.
6 Conclusions
In this paper, we explored a motivated direction for few-shot fake news detection. We began by analyzing the limitations of current LLM-based detection methods, identifying two primary challenges: (1) Understanding Ambiguity and (2) Information Scarcity. To address these issues, we developed a Dual-perspective Knowledge-guided Fake News Detection (DKFND) model. In DKFND, a Detection Module was designed to identify knowledge concepts from given news. Then, we proposed an Investigation Module, a Judge Module to retrieve and select valuable information. Importantly, a Determination Module integrated predictions from both inside and outside perspectives to produce the final output. Finally, extensive experiments demonstrated the effectiveness of our method.
7 Limitation
In our proposed DKFND method, we need to integrate the LLM (i.e., zephyr-7b-beta introduced in Section 5.1). Due to the large scale of LLMs, it tends to consume more computing resources and time compared to traditional baselines, such as PSM (Ni et al., 2020) and FakeFlow (Ghanem et al., 2021). Essentially speaking, LLMs contain a vast amount of knowledge, much of which may be unnecessary for fake news detection. Distilling useful knowledge so as to accelerate the inference remains a valuable and intriguing research direction.
Another limitation is that our current approach only employs LLMs for inference. Although we design precise prompts to implement in-context learning, it still cannot fully exploit the capabilities of LLMs due to the inherent gap between the natural language and the knowledge encoded in the model parameters. This also results in the bad case we analyzed in Section A. In future work, we would like to explore the low resource scenario fine-tuning techniques (e.g., lora (Hu et al., 2021a)) to adapt LLMs for few-shot fake news detection.
Appendix A Bad Case Analysis
@@
(1) For one thing, as introduced in Section 1, DKFND retrieves similar demonstrations from the training set in response to the Understanding Ambiguity problem. This relies on the assumption that valuable samples can be found in the training set, which is not always the case. As a consequence, the failure in Figure 5 occurs. Therefore, how to discriminate and mitigate the impact of such circumstances is a crucial direction for improving the DKFND design. (2) For another, as introduced in Section 4, we only employ the LLM to conduct inference, which cannot fully exploit LLMs’ powerful capabilities. In some cases, although Outside Investigation retrieves valuable information, the prediction of Outside Judge still goes wrong. In fact, despite the great reasoning ability of these general LLMs, they are not competent in news-related domains and are not sufficiently familiar with the specific expression characteristics. Thence, we would like to adopt fine-tuning techniques to adapt LLMs for the news corpus in future, which we believe could bring a positive effect to our model.
Appendix B Outside Investigatation Illustration
For better illustrate the search process of the outside investigation, we give an example in Figure 6. You can also refer to our released anonymous codes for more technical details.
Figure 6: Illustration of the Google search process.
Appendix C Prompts
In this section, we illustrate the prompts utilized in our DKFND methodology, which can serve as a valuable resource for future research in this area.
C.1 Detection Prompt
There is a news article and a list of concepts related to it. Your task is to determine how relevant these keywords are to the given news. Assign each keyword a relevancy score between 0 and 1, where 1 indicates the keyword is highly relevant, and 0 indicates it is not relevant at all."
Instructions:
1. Analyze the content of the news article to understand its main topics and themes.
2. Compare each concept to the news content and assign a relevancy score between 0 and 1.
3. Output the concepts and their scores in reverse order (from the most relevant to the least relevant).
4. Don’t return your explanation about the concepts and their scores.
Output Format:
[[\"concept 1\", <score>], [\"concept 2\", <score>],..., [\"concept N\", <score>]]
Example Input:
News Article:
Scientists have discovered a new method to reduce carbon emissions using advanced nanotechnology. This breakthrough could significantly impact efforts to combat climate change globally.
concepts:
Global warming, Carbon emissions, Cryptocurrency, Nanotechnology, Climate change
Example Output:
[[\"Global warming\", 0.8], [\"Carbon emissions\", 1.0], [\"Cryptocurrency\", 0.0], [\"Nanotechnology\", 1.0], [\"Climate change\", 0.9]]
Input:
News Article:
{Target News Document}.
concepts:
{Target News Concepts}.
Your Output Format:
[[\"concept 1\", #score#], [\"concept 2\", #score#],..., [\"concept N\", #score#]]
Please replace #score# in Your Output Format with the relevancy score of corresponding word:
C.2 Inside Judge Prompt
I need your help determining the reliability of some passages, and I’ve provided some relevant entities mentioned in the text. Here are some hints:
1. **Entity Accuracy**: Check if the entities (e.g., names, attributes, roles) match the passages I’ve provided. Note that we do not require all relevant entities to appear in the passage, you only need to verify that the entities present in the passage do not conflict with the entities provided. You also do not need to check the relationships between the entities. However, if none of the entities appear, then the passage should be unreliable.
Please output your reason and index of the two most relevant passage in the following JSON format:
{"reason": <your explanation for the decision>,"index": [<index of passage which is the most relevant>, <index of passage which is the second relevant>]}
Here are the Passages:
{Retrieved News Documents}.
Here are the relevant entities:
{Current News Concepts}.
Output:
{"reason": #Explanation#,"index": [#Index 1#, #Index 2#]}
Please replace #Explanation# with your explanation for the decision and replace #Index 1#, #Index 2# with index of passage which is the first relevant and second relevant.
C.3 Outside Judge Prompt
## Task Name:
Extract Relationships between Provided Entities from a Text.
## Task Description:
Your task is to extract triples that represent relationships between a set of **predefined entities** based on the provided text. The subject and object of each triple must come **only** from the list of entities that I provide. If no meaningful relationships are found between the entities, return **None**.
### Task Input:
1. A piece of text.
2. A list of possible entities.
### Task Output:
A list of triples in JSON format, or **None** if no valid relationships are found.
### JSON Format for Triples:
[{"subject": "subject_name", "predicate": "predicate_relation", "object": "object_name"}, {"subject": "subject_name","predicate": "predicate_relation","object": "object_name"}]
### Guidelines:
- The **subject** and **object** of the triple must come from the list of predefined entities. Do not extract any other words as entities.
- If there are no valid relationships between the provided entities, return **None**.
### Examples:
#### Example 1:
**Input Text:**
"Paris is the capital of France."
**Entities Provided:**
["Paris", "France"]
**Output JSON:**
[{"subject": "Paris", "predicate": "is the capital of", "object": "France"}]
#### Example 2:
**Input Text:**
"Einstein developed the theory of relativity."
**Entities Provided:**
["Einstein", "the theory of relativity"]
**Output JSON:**
["subject": "Einstein", "predicate": "developed", "object": "the theory of relativity"]
#### Example 3:
**Input Text:**
"She is a member of the organization."
**Entities Provided:**
["the organization"]
**Output JSON:**
None
#### Example 4:
**Input Text:**
"The Nile is the longest river in the world, flowing through Egypt."
**Entities Provided:**
["The Nile", "Egypt"]
**Output JSON:**
[{"subject": "The Nile", "predicate": "flows through", "object": "Egypt"}]
### Real Data:
**Input Text:**
{Target News Document}.
**Entities Provided:**
{Target News Concepts}.
**Output JSON:**
[{"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}, {"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}, {"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}]
Please replace ##predefined_entity## in **Output JSON** of Real Data with entities from Entities Provided and replace #Relation# in **Output JSON** of Real Data with meaningful relationship between entities.
I do not need you give me code, return triples in JSON format, or **None** if no valid relationships are found.
C.4 Inside Determination Prompt
You are an expert of news authenticity evaluation. As an expert of news authenticity evaluation, you should analyze and evaluate the authenticity of news.
I need your assistance in evaluating the authenticity of a news article. Please focus on whether the content of the article is true, not on the genre and rhetoric of the article.
I will provide you the news article and additional information about this news.
You have to answer that [This is fake news] or [This is real news] in the first sentence of your output and give your explanation about [target news].
[example i]:
[input news]: [{}]
[output]: [This is real/fake news]
[target news]: [{}]
[input news]: [{}]
[output]:
C.5 Outside Determination Prompt
I need your assistance in evaluating the authenticity of a news article. I will provide you the news article and additional information about this news. Please analyze the following news and give your decision and reason. The first sentence of your [Decision and Reason] must be [This is fake news] or [This is real news], and then give reason.
The news article is:
{Target News Document}.
The additional information is:
{Outside Information}.
[Decision and Reason]:
C.6 Integrated Determination Prompt
I need your assistance in evaluating the authenticity of a news article. This news article include news title, news text and news tweet.
Here are the news article:
{Target News Document}.
There are two different views on this news article.
Some people believe that {}, their explanation is: {}
Others believe that {}, their explanation is: {}
Please judge their opinion, evaluate the authenticity of the news article by analyzing its logic, reliability, and factual accuracy, and finally give your decision.
The first sentence after [Explanation] must be [This is fake news] or [This is real news], and then give your explanation.
[Explanation]:
C.7 Simplified Integrated Determination Prompt for Ablation Study
I need your assistance in evaluating the authenticity of a news article. This news article include news title, news text and news tweet.
Here are the news article:
{Target News Document}.
There are two different views on this news article.
Some people believe that {}, their explanation is: {}
Others believe that {}, their explanation is: {}
Please give your decision.
The first sentence after [Explanation] must be [This is fake news] or [This is real news], and then give your explanation.
[Explanation]:

### v5

@@ V-B Experimental Result
The main results, presented in Table II, indicate that our proposed DKFND model surpasses all baselines across various metrics, encompassing traditional, LLM-based and hybrid methods. This underscores the effectiveness of our design and the advantages of enhancing the LLM through both inside and outside perspectives. Furthermore, several notable phenomena emerge from these results:
@@
Firstly, for most baselines and our DKFND model, the performance on the PolitiFact dataset exceeds that on the Gossipcop dataset, suggesting that Gossipcop presents greater difficulty. Specifically, PolitiFact consists of political news while Gossipcop pertains to the entertainment domain. This disparity is reasonable as political news typically exhibits more organized format and content, which facilitates the fake news detection process. Secondly, with the increase of training instances ($K$), most traditional fake news detection methods (e.g., ④ PSM) and hybrid methods ( 11 ARG) show improved performance. This is logical as more data enables better training of a supervised model, mitigating the lack of prior knowledge. However, an exception is observed in ① PROPANEWS, whose performance appears relatively unaffected by $K$. As introduced in Section V-A2, different from other methods, PROPANEWS designs a data constructing strategy to supplement original training set. This significantly offsets the impact of training data quantity. Moreover, for most LLM-based methods (⑦ $\sim$ 10), as outlined in Section V-A2, due to the limitation of maximum tokens, we all randomly select 6 samples as demonstrations. Hence, increasing the number of training instances does not substantially benefit the in-context learning of LLMs. Thirdly, although our DKFND model also faces the constraint of maximum token limitations, the $k$NN retrieval mechanism and inside judge designs enhance the utilization of increased training data, thereby achieving a certain degree of improvements with higher $K$ values. Fourthly, the hybrid method ( 11 ARG), benefiting from the joint modeling of traditional models and LLMs, obtains competitive performance. And compared to that, DKFND still maintains a significant advantage, particularly in scenarios with scarcer data. These observations further demonstrate the effectiveness of our designs from multiple perspectives.
V-C Component Effectiveness Analysis
@@
In this subsection, we conduct ablation experiments to assess the effectiveness of different components within our model. Specifically, we first simplify the Integrated Determination part (Section IV-D1), removing the detailed explanation in promot $P_{FD}$ (Eq.(10)). On the basis of this, we design ablated variants from inside and outside perspectives.
@@
  • DKFND-Inside: simplifying the design of Section IV-B1 and IV-C1 by randomly selecting samples from training set as demonstrations.
@@
  • DKFND-Outside: simplifying the design of Section IV-C2 by randomly selecting retrieved outside information.
@@
Figure 4: The case study of the DKFND model. They come from the PolitiFact dataset ($K$=100).
@@
Furthermore, considering that the backbone of DKFND is ⑦ Zephyr[17], we can jointly compare the results in Figure 3 and the “⑦ Zephyr” line in Table II. This comparison further reveals that all three ablation variants outperform Zephyr, demonstrating that even in its simplified forms, leveraging both internal and external perspectives for fake news detection yields significant performance gains. This finding further validates the motivation behind our dual-perspective design.
V-D Case Study
@@
VI Bad Case Analysis
@@
(1) For one thing, as introduced in Section I, DKFND retrieves similar demonstrations from the training set in response to the Understanding Ambiguity problem. This relies on the assumption that valuable samples can be found in the training set, which is not always the case. As a consequence, the failure in Figure 5 occurs. Therefore, how to discriminate and mitigate the impact of such circumstances is a crucial direction for improving the DKFND design. (2) For another, as introduced in Section IV, we only employ the LLM to conduct inference, which cannot fully exploit LLMs’ powerful capabilities. In some cases, although Outside Investigation retrieves valuable information, the prediction of Outside Judge still goes wrong. In fact, despite the great reasoning ability of these general LLMs, they are not competent in news-related domains and are not sufficiently familiar with the specific expression characteristics. Therefore, we would like to adopt the fine-tuning techniques to adapt LLMs for the news corpus in future, which we believe could bring a positive effect to our DKFND model.
VII Limitation
In our proposed DKFND method, we need to integrate the LLM (i.e., zephyr-7b-beta introduced in Section V-A). Due to the large scale of LLMs, it tends to consume more computing resources and time compared to traditional baselines, such as PSM[36] and FakeFlow[10]. Essentially speaking, LLMs contain a vast amount of knowledge, much of which may be unnecessary for fake news detection. Distilling useful knowledge so as to accelerate the inference remains a valuable and intriguing research direction.
Another limitation is that our current approach only employs LLMs for inference. Although we design precise prompts to implement in-context learning, it still cannot fully exploit the capabilities of LLMs due to the inherent gap between the natural language and the knowledge encoded in the model parameters. This also results in the bad case we analyzed in Section VI. In future work, we would like to explore the low resource scenario fine-tuning techniques (e.g., lora[41]) to adapt LLMs for few-shot fake news detection.
VIII Conclusions
In this paper, we explored a motivated direction for few-shot fake news detection. We began by analyzing the limitations of current LLM-based detection methods, identifying two primary challenges: (1) Understanding Ambiguity and (2) Information Scarcity. To address these issues, we developed a Dual-perspective Knowledge-guided Fake News Detection (DKFND) model. In DKFND, a Detection Module was designed to identify knowledge concepts from given news. Then, we proposed an Investigation Module, a Judge Module to retrieve and select valuable information. Importantly, a Determination Module integrated predictions from both inside and outside perspectives to produce the final output. Finally, extensive experiments on two real-world datasets demonstrated the effectiveness of our method.